OpenStudy (cpinder):

Estimate the temperature at which 40 kg of ethane stored in a 0.15 m3 vessel exerts a pressure of 20000 kPa.

OpenStudy (cpinder):

So far, I've got calculated the volume like this. Vt=0.15m3=150000cm3 V=V/MW=150000cm3/(40kg*30mol/kg) = 112500 cm3/mol And I've calculated Pr like so. Pr=P/Pc= (20000 kPa*.01 bar/kPa) / 48.72 bar = 4.105 I'm lost from here.

OpenStudy (cpinder):

I'm looking at the solution manual right now and this is what it says. PV=ZRT=ZRTrTc Tr=alpha/Z where alpha=PVR/Tc alpha=29.548mol/kg (Tried to get this but I couldn't) whence Tr=0.889/Z at Pr=P/Pc (I've calculated Pr already) This equation giving Tr as a function of Z and eq (3.57) in conjunction with tables E3 and E4 are two relations in the same variables which must be satisfied at the given reduced pressure. The intersection of these two relations can be found by one means or another to occur at about: Tr=1.283 and Z=0.693 T=Tr*Tc=391.7K or 118.5C

OpenStudy (cpinder):

@surry99 any idea?